# Twist, splay, and uniform domains in ferroelectric nematic liquid crystals

### v1

@@ III.1 Twist in a cylinder
We find that, given the reasonable parameters mentioned above and $T=350~{}\mathrm{K}$, $c^{*}\approx 3\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$, which is on the same order of magnitude as the concentration of ions measured in the N phase, $c=(5-6)\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$, at which the TD domains are observed, Fig. 1(a, c-f). However, the experimental data obtained in the N phase might not be representative of the concentration of ions in the $\mathrm{N}_{\mathrm{F}}$ phase. Nevertheless, the theoretical estimate appears to be consistent with the experiments in which the TDs disappear when DIO is doped with the ionic fluid BMIM-$\mathrm{PF}_{\mathrm{6}}$ [see Fig. 1(b)], which could potentially increase the concentration of ions by orders of magnitude.
III.2 $\pi$-twists
We have so far considered thick cells with the polarization $\mathbf{P}$ twisting in a cylinder along the direction normal to the cell (the $z$-axis). In this case, the elastic and electrostatic energies compete throughout the bulk of the sample. However, we now consider an $\mathrm{N}_{\mathrm{F}}$ confined to a thin cell with boundary conditions, such as a certain imposed nematic orientation $\mathbf{n}_{0}$ at the top and bottom of the cell, as in the samples shown in Fig. 1. When the anchoring is strong, the sample thickness $h$ will typically set the periodicity of the twist along the $z$ direction. This is demonstrated by polarizing light microscopy (see Basnet et al. (2022) for details) in Fig. 1(c-f), which shows that most domains in the cells exhibit a $\pi$-twist in the polarization $\mathbf{P}$ from the bottom to the top surface.
However, for even thinner cells, the $\pi$-twist becomes too energetically costly. Indeed, there is a critical thickness $h^{*}$ for which we get twisted domains if $h>h^{*}$ and domains with uniform $\mathbf{P}$ orientation for $h<h^{*}$. We can calculate the critical thickness $h^{*}$ by considering a simple square domain with dimension $L$ in the cell of thickness $t$. At sufficiently large anchoring strengths, we may assume that $\mathbf{P}$ remains in the $xy$ plane and that $\mathbf{P}$ runs along $\hat{\mathbf{y}}$ on the top and bottom surfaces. We compare two polarization configurations: a uniform $\mathbf{P}_{\mathrm{uniform}}=P_{0}\hat{\mathbf{y}}$ and a $\pi$-twisted $\mathbf{P}_{\pi-\mathrm{twist}}=P_{0}\sin(\pi z/h)\hat{\mathbf{x}}+P_{0}\cos(\pi z/h)\hat{\mathbf{y}}$, with both vanishing outside of the region $-L/2<x,y<L/2$ and $0<z<h$. The Fourier transforms are
$\begin{cases}\tilde{\mathbf{P}}_{\mathrm{uniform}}=\frac{8P_{0}e^{-ihk_{z}/2}\sin\left(\frac{k_{z}h}{2}\right)\sin\left(\frac{k_{x}L}{2}\right)\sin\left(\frac{k_{y}L}{2}\right)\hat{\mathbf{y}}}{k_{x}k_{y}k_{z}}\\ \tilde{\mathbf{P}}_{\pi-\mathrm{twist}}=\frac{8P_{0}\sin\left(\frac{k_{x}L}{2}\right)\sin\left(\frac{k_{y}L}{2}\right)\cos\left(\frac{hk_{z}}{2}\right)[\pi\hat{\mathbf{x}}+ihk_{z}\hat{\mathbf{y}}]}{k_{x}k_{y}[\pi^{2}-h^{2}k_{z}^{2}]e^{\frac{ihk_{z}}{2}}}\,\frac{}{}\end{cases}.$ (14)
Substituting these transforms into Eq. (2) and assuming that $L\kappa\gg 1$ (strong screening or large sample size) yields the dipolar energy
@@ III.2 $\pi$-twists
The $\pi$-twisted configuration incurs an elastic energy penalty given by $F_{n}=K_{2}L^{2}\pi^{2}/(2h)$, which follows from substituting $\mathbf{n}_{\mathrm{twist}}=\sin(\pi z/h)\hat{\mathbf{x}}+\cos(\pi z/h)\hat{\mathbf{y}}$ into the Frank free energy, Eq. (3). The elastic energy decreases with cell thickness $h$, while the dipolar energy in Eq. (15) increases. The balance yields a critical thickness
$h^{*}=\sqrt{\frac{2\epsilon\epsilon_{0}\kappa K_{2}L\pi^{2}}{3P_{0}^{2}}}.$ (16)
The critical thickness $h^{*}$ increases with the domain size $L$ because the dipolar energy comes from uncompensated charge at the boundary of the domain. However, even for extremely large domains with $L\approx 1~{}\mathrm{mm}$, we find a remarkably small $h^{*}\approx 1-10~{}\mu\mathrm{m}$ for the experimental conditions considered here (assuming $\lambda_{D}=\kappa^{-1}\approx 0.1-10~{}\mu\mathrm{m}$). In other words, the depolarization field plays a significant role even though it is essentially a boundary effect. We expect to see $\pi$-twisted domains, such as those shown in Fig. 1(a,c-f), even for cells with micron-scale thickness $h$. In the experiments, we find that essentially all domains are twisted for thicknesses $h>2~{}\mu\mathrm{m}$, consistent with this result.
Confinement can induce chirality (twist) in solid state ferroelectrics, as well, especially in nanostructured materials Lukyanchuk et al. (2024, 2025), although intrinsically chiral solid ferroelectrics are also possible. As we shall see in the next section, the polarization can also be modulated in the perpendicular direction (along the cell boundary). Such modulated states exist in the solid state Lukyanchuk et al. (2025), even with smooth variation of the polarization direction. We will thus draw some additional comparisons between the crystalline and nematic ferroelectrics in the following.
IV Domain patterns
Apart from the $\pi$-twist along the $z$ direction, we know that the $\mathrm{N}_{\mathrm{F}}$ in these thin cells breaks up into domains along the cell surface. We will study some prototypical cases, including the stripes in Fig. 1 and the pie slices in Figs. 2,3. Let us begin with an analysis of the stripes in Fig. 1.
IV.1 Stripe domains
@@
Figure 6: (a) Schematic of a possible polarization direction in the striped domains with characteristic wavelength $\lambda_{x}$, with the anchoring nematic direction $\mathbf{n}_{0}$ on the top and bottom surfaces. (b) Dimensionless free energy [Eq. (20)] of the striped domain configuration, taking into account both the electrostatic energy contribution and the energy of domain walls. We see that there is a minimum at a certain $\kappa\lambda_{x}$, indicating that the striped configuration is favorable. We show the energy for various values of the dimensionless parameter $\kappa\lambda_{\mathrm{dw}}\sqrt{L/h}$ discussed in the text.
@@
The domains observed in Fig. 1 are reminiscent of structures found in thin films of solid, uniaxial ferroic materials Fernandez et al. (2022), where the polarization $\mathbf{P}$ (or magnetization for ferromagnets) forms stripes of alternating orientation (although the polarization typically has components perpendicular to the film surface, unlike the $\mathrm{N}_{\mathrm{F}}$ which has $\mathbf{P}$ parallel to the $xy$ plane) Streiffer et al. (2002); Feigl et al. (2014). Like the domains in the $\mathrm{N}_{\mathrm{F}}$ and the cylindrical case considered above (see Fig. 5), the stripes in solid ferroics are generated to mitigate the strong depolarization field Kittel (1946); Stephenson and Elder (2006): A thick stripe will break up into thinner stripes in order to create more cancellation of the uncompensated charge at the stripe boundaries. Similar striped patterns appear in ferromagnetic crystals, with the patterns analyzed 90 years ago by Landau and Lifshits Landau and Lifshits (1935).
@@
Consider a thin, square cell with thickness $h$ and square cross section with area $A_{xy}=L^{2}$, where $L=L_{x}=L_{y}$ is the linear extent of our sample. The basic idea, also discussed in detail by Kittel for solid ferromagnetic materials Kittel (1949), is that the dipolar energy density $f_{\rho}=F_{\rho}/A_{xy}$ per area of the cell will scale approximately as $f_{\rho}\propto\lambda_{x}$, with $\lambda_{x}$ the stripe size for stripes running along the $y$-axis, say. Meanwhile, the introduction of alternating domains of $\mathbf{P}$ will incur a cost due to the domain walls. There will be approximately $2L/\lambda_{x}$ domain walls in the sample, so the associated free energy density will scale according to $f_{\mathrm{dw}}\propto 1/\lambda_{x}$. We see, therefore, that there should be an optimal value of $\lambda_{x}$ which balances the two energies $f_{\mathrm{dw}}$ and $f_{\rho}$.
@@
Let us analyze the optimal wavelength $\lambda_{x}$ by assuming that we have very strong anchoring so that within each stripe we have a uniform $\pi$-twisted polarization $\mathbf{P}=\pm\mathbf{P}_{\pi-\mathrm{twist}}$ which satisfies the bidirectional boundary conditions at the cell surface, as illustrated in Fig. 6(a). Between each stripe domain at the cell surfaces, we have a 180 flip in the polarization orientation. One possibility, illustrated in Fig. 6(a), is that adjacent domains are of opposite chirality, so that the polarization field is uniform in the middle of the cell at $z=h/2$. The corresponding polarization field $\mathbf{P}_{\mathrm{stripe}}$ for this configuration reads
$\mathbf{P}_{\mathrm{stripe}}$ $=\frac{4P_{0}}{\pi}\,\sum_{n=1}^{\infty}\,\frac{1}{n}\,\sin\left(\frac{\pi n}{2}\right)\cos(nq_{x}x)\cos\left(\frac{\pi z}{h}\right)\hat{\mathbf{y}}$
${}+P_{0}\sin\left(\frac{\pi z}{h}\right)\hat{\mathbf{x}},$ (17)
@@
where $q_{x}=2\pi/\lambda_{x}$ is the wavevector associated with the stripe wavelength $\lambda_{x}$ [see red double arrow in Fig. 6(a)]. The summation over $n$ is the mode expansion of a square wave, so that we have a rapid reorientation of $\mathbf{P}$ from $+P_{0}\hat{\mathbf{y}}$ to $-P_{0}\hat{\mathbf{y}}$ at the top and bottom of the cell ($z=0,h$), as illustrated in Fig. 6(a). Let us assume that the striped configuration has some lateral extent $0<x,y<L$, respectively. For simplicity, we also assume that the total sample is neutral, which means that $L$ should be some integer multiple of $\lambda_{x}$.
We substitute the Fourier-transformed Eq. (17) into Eq. (2) to find the dipolar energy of this configuration. Much like the cylinder domain considered previously, we expect that the electrostatic energy will increase with increasing $\lambda_{x}$, as larger $\lambda_{x}$ means that we have more uncompensated charge at the boundaries of the stripes. We find, for large sample sizes $L\kappa\gg 1$, that
@@
$F_{\rho}$ $=\frac{4P_{0}^{2}}{\pi^{3}\epsilon\epsilon_{0}}\,\iiint_{-\infty}^{\infty}\frac{\cos^{2}\left(\frac{z}{2}\right)\sin^{2}\left(\frac{xL}{2h}\right)\sin^{2}\left(\frac{yL}{2h}\right)}{(x^{2}+y^{2}+z^{2}+h^{2}\kappa^{2})(\pi^{2}-z^{2})^{2}}\left[\frac{16x^{2}z^{2}}{\pi^{2}h}\left[\sum_{n=1}^{\infty}\frac{h^{2}\lambda_{x}^{2}\sin\left(\frac{\pi n}{2}\right)}{n(\lambda_{x}^{2}x^{2}-4\pi^{2}n^{2}h^{2})}\right]^{2}+\frac{h^{3}\pi^{2}}{y^{2}}\right]\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}z$ (18)
@@
As long as the screening contribution in the sum in Eq. (19) is negligible, then the electrostatic energy $F_{\rho}$ grows with $\lambda_{x}$. (We expect this to happen whenever the equilibrium $\lambda_{x}$ satisfies $\lambda_{x}\kappa\lesssim 1$.) So, the electrostatic energy prefers to have a small $\lambda_{x}$. However, decreasing $\lambda_{x}$ will introduce more domain walls into the system.
The nature of the domain walls may be complex due to the twist in the polarization $\mathbf{P}$. Experiments indicate that the domain wall may consist of surface disclination lines. These lines come in pairs, as the polarization is uniform at $z=h/2$ Chen et al. (2020); Yi et al. (2024); Kumari et al. (2024b). An estimate for the energy of the wall (per unit length) would be $f^{\mathrm{disc}}_{\pi}\approx 2K$, with $K$ an elastic constant. Another possibility is that the domain wall is a solitonic, Bloch wall structure, which we may call a $\pi$-wall as $\mathbf{P}$ flips by 180 across the wall Basnet et al. (2022). The energy per unit length of one such $\pi$-wall is given by $f^{\mathrm{wall}}_{\pi}=2\sqrt{2KhW}$, where $K$ is an elastic constant, $W$ is the anchoring strength, and $h$ is the cell thickness. The total elastic cost of domain walls is thus approximately $F_{\mathrm{dw}}\approx 2f_{\pi}L^{2}/\lambda_{x}$, with $f_{\pi}$ either the solitonic wall $(f_{\pi}^{\mathrm{wall}})$ or disclination line pair $(f_{\pi}^{\mathrm{disc}})$ energy density.
@@
$\frac{F}{F_{0}}$ $=\frac{1}{4}+\frac{\kappa\lambda_{x}}{\pi^{2}}\sum_{n=1}^{\infty}\,\frac{[1-(-1)^{n}]}{n^{2}\sqrt{(2\pi n)^{2}+(\kappa\lambda_{x})^{2}}}+\,\frac{(\kappa\lambda_{\mathrm{dw}})^{2}L}{\kappa\lambda_{x}h}$
$\approx\frac{1}{4}+\frac{2\kappa\lambda_{x}}{\pi^{2}\sqrt{4\pi^{2}+(\kappa\lambda_{x})^{2}}}+\,\frac{(\kappa\lambda_{\mathrm{dw}})^{2}L_{y}}{\kappa\lambda_{x}h},$ (20)
where $F_{0}=hP_{0}^{2}L(\epsilon\epsilon_{0}\kappa)^{-1}$ is a characteristic free energy and $\lambda_{\mathrm{dw}}=(2f_{\pi}\epsilon\epsilon_{0})^{1/2}P_{0}^{-1}$ is a characteristic length associated with the domain wall. Note that the length $\lambda_{\mathrm{dw}}$ is quite small (on the order of nanometers) and arises from the interplay between the elastic cost of domain walls and the electrostatics. A similar length was derived some time ago by balancing analogous energetic contributions Zhuang et al. (1989). We have made an additional assumption that we can take $h\kappa\gg 1$ in the left-over integration in Eq. (19), which will not change the location of the minimum of $F$ with respect to $\lambda_{x}$, which we will now explore.
We plot the dimensionless free energy $F/F_{0}$ in Eq. (20) versus $\kappa\lambda_{x}$ in Fig. 6(b) for various values of $\kappa\lambda_{\mathrm{dw}}\sqrt{L/h}$, with $L$ the linear extent of the sample. It is possible to numerically minimize the function in Eq. (20), which demonstrates that for $\kappa\lambda_{\mathrm{dw}}\sqrt{L/h}<0.618$, the free energy curve has a global minimum at a finite value of $\lambda_{x}$, indicating that the system prefers to make stripes of a certain size. For larger values, the free energy minimum corresponds to $\lambda_{x}\rightarrow\infty$ and a uniform polarization state. In the low screening limit $\kappa\lambda_{\mathrm{dw}}\ll\sqrt{h/L}$, the preferred wavelength (free energy minimum) is at, approximately,
@@
IV.2 Pie slice domains
Figure 7: (a) Schematic of the polarization orientation near the cell surface for a cell pre-patterned with a $+1$ defect. Apart from the domain walls (along the $x$ and $y$ axes and along the dashed line), the polarization $\mathbf{P}$ runs along the radial direction $\hat{\mathbf{r}}$. The polarization configuration has $n_{\theta}=4$ pairs of “pie-slice” sectors of anti-parallel $\mathbf{P}$. (b) Plot of the rescaled free energy $F/F_{0}$ in Eq. (27) of a pie-slice pattern with $n_{\theta}$ cuts for various values of $h\kappa$, with $\kappa$ the inverse Debye screening length and $h$ the cell thickness. We fix $R/h=100$ and $\eta_{0}=30$. We see that the Coulomb and elastic energies balance to generate an optimal value of $n_{\theta}$ (the free energy minimum). The dashed vertical line indicates the optimal value in the low screening $\kappa\rightarrow 0$ limit.
We now consider the $+1$ aster defect anchoring condition, with $\mathbf{n}_{0}=\hat{\mathbf{r}}$ the preferred orientation at the top and bottom cell surfaces. A pure $+1$ aster defect in the polarization vector, $\mathbf{P}=P_{0}\hat{\mathbf{r}}$, has a corresponding charge density distribution
@@ IV.2 Pie slice domains
Note that the charge is no longer confined to the boundary, as in the situations considered previously (the cylindrical and striped domains). Instead, there is a distribution of bound charge concentrated at the origin of the defect. Like the boundary charges, this charge is energetically costly, so the $\mathrm{N}_{\mathrm{F}}$ prefers to reorient along the $\hat{\bm{\theta}}$ direction to create a net neutral charge configuration, forming a series of $n_{\theta}$ pairs of alternating $\mathbf{P}$ domains (with $\mathbf{P}\parallel\pm\hat{\mathbf{r}}$), as shown in Fig. 7(a). Of course, we also expect a $\pi$-twist along the $z$-direction, but we will neglect the $z$-dependence for simplicity, focusing on the regions of the $\mathrm{N}_{\mathrm{F}}$ near the cell surface, where $\mathbf{P}$ is forced parallel to the radial direction due to anchoring conditions.
So, consider the polarization $\mathbf{P}$ configuration in a circular domain of radius $R$, confined to a cell with thickness $h$. Far from the defect, as $R\rightarrow\infty$, we expect to see striped domains with the preferred wavelength $\lambda_{x}^{*}$ calculated in the previous section. However, near the defect, we expect the bound charge distribution in the bulk should play a more significant role. We see in experiments that the sample breaks up into “pie slice” domains with sharp tips, as shown in Fig. 7(a,b). Assuming the polarization remains in the $xy$ plane and does not depend on the distance $r$ from the defect, the form of the polarization vector of such a configuration is

### v2

@@ Results
Given the reasonable parameters mentioned above and $T=350~{}\mathrm{K}$, we find $c^{*}\approx 5\times 10^{21}~{}\mathrm{ions}/\mathrm{m}^{3}$, which is less than the concentration of ions measured in the N phase, $c=(5-6)\times 10^{22}~{}\mathrm{ions}/\mathrm{m}^{3}$. However, the experimental data obtained in the N phase might not be representative of the concentration of ions in the $\mathrm{N}_{\mathrm{F}}$ phase. Also, our analysis so far has not taken into account the cell boundary conditions. Nevertheless, the theoretical estimate appears to be consistent with the experiments in which the TDs disappear when DIO is doped with the ionic fluid BMIM-$\mathrm{PF}_{\mathrm{6}}$ [see Fig. 1f], which could increase the concentration of ions by orders of magnitude.
Critical cell thickness for $\pi$-twists
We now consider an $\mathrm{N}_{\mathrm{F}}$ confined to a cell with certain imposed nematic orientation $\mathbf{n}_{0}$ at the top and bottom of the cell, as in the samples shown in Fig. 1 with strong anchoring. There is a critical thickness $h^{*}$ for which we get twisted domains if $h>h^{*}$ and domains with uniform $\mathbf{P}$ orientation for $h<h^{*}$. When $h>h^{*}$, the sample thickness $h$ will set the periodicity of the twist along the $z$ direction (see Fig. 1e-j).
We can calculate the critical thickness $h^{*}$ by considering a square domain with dimension $L$. We assume that $\mathbf{P}$ remains in the $xy$ plane and that $\mathbf{P}$ runs along $\hat{\mathbf{y}}$ on the top and bottom surfaces. We compare two polarization configurations: a uniform $\mathbf{P}_{\mathrm{uniform}}=P_{0}\hat{\mathbf{y}}$ and a $\pi$-twisted $\mathbf{P}_{\pi-\mathrm{twist}}=P_{0}\sin(\pi z/h)\hat{\mathbf{x}}+P_{0}\cos(\pi z/h)\hat{\mathbf{y}}$, with all polarizations vanishing outside of the region $-L/2<x,y<L/2$ and $0<z<h$. The Fourier transforms are
$\begin{cases}\tilde{\mathbf{P}}_{\mathrm{uniform}}=\dfrac{8P_{0}\sin\left(\frac{k_{z}h}{2}\right)\sin\left(\frac{k_{x}L}{2}\right)\sin\left(\frac{k_{y}L}{2}\right)\hat{\mathbf{y}}}{k_{x}k_{y}k_{z}e^{\frac{ihk_{z}}{2}}}\\[8.0pt] \tilde{\mathbf{P}}_{\pi-\mathrm{twist}}=\dfrac{8P_{0}h\sin\left(\frac{k_{x}L}{2}\right)\sin\left(\frac{k_{y}L}{2}\right)\cos\left(\frac{hk_{z}}{2}\right)[\pi\hat{\mathbf{x}}+ihk_{z}\hat{\mathbf{y}}]}{k_{x}k_{y}[\pi^{2}-h^{2}k_{z}^{2}]e^{\frac{ihk_{z}}{2}}}\end{cases}.$ (14)
Substituting Eq. (14) into Eq. (2) and assuming that $L\kappa\gg 1$ (strong screening or large sample size) yields the dipolar energy
@@
The $\pi$-twisted configuration incurs an elastic energy penalty given by $F_{n}=K_{2}L^{2}\pi^{2}/(2h)$, which follows from substituting $\mathbf{n}_{\pi-\mathrm{twist}}=\sin(\pi z/h)\hat{\mathbf{x}}+\cos(\pi z/h)\hat{\mathbf{y}}$ into the Frank free energy, Eq. (3). The balance between elastic and electrostatic energies yields a critical thickness
$h^{*}=\frac{\pi}{P_{0}}\sqrt{\frac{2\epsilon\epsilon_{0}\kappa K_{2}L}{3}}.$ (16)
Even for large domains with $L\approx 1~{}\mathrm{cm}$, we find a small $h^{*}\approx 0.2-2~{}\mu\mathrm{m}$, for a wide range of screening lengths $\lambda_{D}=\kappa^{-1}\approx 0.01-1~{}\mu\mathrm{m}$. In the experiments, we find that essentially all domains are twisted for thicknesses $h>2~{}\mu\mathrm{m}$, consistent with this result. Confinement can induce chirality (twist) in solid state ferroelectrics, as well, especially in nanostructured materials [25, 26], although intrinsically chiral solid ferroelectrics are also possible.
Model of stripe domain patterns
@@
Figure 6: Striped domain model. a Schematic of a possible polarization direction in the striped domains with characteristic wavelength $\lambda_{x}$, with the anchoring nematic direction $\mathbf{n}_{0}$ indicated on the top and bottom surfaces. The gray surfaces separating the blue and green domains could be either disclination lines, solitons, of fixed-width domain walls spanning the entire cell height $h$. b Dimensionless free energy [Eq. (19)] of the striped domain configuration, taking into account both the electrostatic energy contribution and the energy of domain walls (captured by the parameter $\lambda_{\mathrm{dw}}$). We see that there is a minimum at a certain $\kappa\lambda_{x}$, indicating that the striped configuration is favorable. We show the energy for various values of the dimensionless parameter $\kappa\lambda_{\mathrm{dw}}\sqrt{L/h}$ discussed in the text.
@@
The domains in Fig. 1 are reminiscent of structures found in thin films of solid, uniaxial ferroic materials [27], where the polarization $\mathbf{P}$ forms stripes of alternating orientation (although the polarization typically has components perpendicular to the film surface, unlike the $\mathrm{N}_{\mathrm{F}}$ which has $\mathbf{P}$ parallel to the $xy$ plane) [28, 29]. The stripes in solid ferroics are generated to mitigate the depolarization field [30, 31]. Similar striped patterns appear in ferromagnetic crystals, with the patterns analyzed 90 years ago by Landau and Lifshitz [32].
@@
Consider a thin, square cell with thickness $h$ and square cross section with area $A_{xy}=L^{2}$, where $L=L_{x}=L_{y}$ is the linear extent of our sample. The basic idea, also discussed in detail by Kittel for solid ferromagnetic materials [33], is that the dipolar energy density $f_{\rho}=F_{\rho}/A_{xy}$ per area of the cell will scale approximately as $f_{\rho}\propto\lambda_{x}$, with $\lambda_{x}$ the stripe size for stripes running along the $y$-axis, say. Meanwhile, the introduction of alternating domains of $\mathbf{P}$ will incur a cost due to the domain walls. There will be approximately $2L/\lambda_{x}$ domain walls in the sample, so the associated free energy density will scale according to $f_{\mathrm{dw}}\propto 1/\lambda_{x}$. There should be an optimal value of $\lambda_{x}$ which balances the two energy densities $f_{\mathrm{dw}}$ and $f_{\rho}$.
@@
Let us analyze the optimal wavelength $\lambda_{x}$ by assuming that we have very strong anchoring so that each stripe satisfies the bidirectional boundary conditions at the cell surface, as illustrated in Fig. 6a. One possibility, illustrated in Fig. 6a, is that adjacent domains are of opposite chirality. The corresponding polarization field $\mathbf{P}_{\mathrm{stripe}}$ for such a configuration reads
$\mathbf{P}_{\mathrm{stripe}}=\frac{4P_{0}}{\pi}\,\sum_{n=1}^{\infty}\,\frac{1}{n}\,\sin\left(\frac{\pi n}{2}\right)\cos(nq_{x}x)\cos\left(\frac{\pi z}{h}\right)\hat{\mathbf{y}}+P_{0}\sin\left(\frac{\pi z}{h}\right)\hat{\mathbf{x}},$ (17)
@@
where $q_{x}=2\pi/\lambda_{x}$ is the wavevector associated with the stripe wavelength $\lambda_{x}$ (see red double arrow in Fig. 6a). The summation over $n$ is the mode expansion of a square wave, so that we have a rapid reorientation of $\mathbf{P}$ from $+P_{0}\hat{\mathbf{y}}$ to $-P_{0}\hat{\mathbf{y}}$ at the top and bottom of the cell ($z=0,h$), as illustrated in Fig. 6a.
A sudden jump in the polarization (±$\pm P_{0}\hat{\mathbf{y}}$ in Fig. 6a) is unrealistic and the detailed structure near the flip (gray surfaces separating the green and blue arrows in Fig. 6a) could take a variety of forms including disclination line pairs, a solitonic structure, or a fixed-width domain wall, as discussed in more detail below. We assume here that this region does not significantly influence the electrostatic contribution to the energy, instead contributing to the elastic cost of the polarization domain. We also assume that the striped configuration has some lateral extent $0<x,y<L$, where $L$ is some integer multiple of $\lambda_{x}$, for simplicity. Although the system is overall charge neutral, there will be regions of charge at the boundary of the domains along the $x,y$ directions where the polarization arrows terminate, creating depolarization fields.
We substitute the Fourier-transformed Eq. (17) into Eq. (2) to find the dipolar energy of this configuration, again assuming a screened Coulomb interaction everywhere with a screening length $\kappa^{-1}$. We find, for large sample sizes $L\kappa\gg 1$, that
@@
$F_{\rho}$ $=\frac{4P_{0}^{2}}{\pi^{3}\epsilon\epsilon_{0}}\,\int\frac{\cos^{2}\left(\frac{z}{2}\right)\sin^{2}\left(\frac{xL}{2h}\right)\sin^{2}\left(\frac{yL}{2h}\right)}{(x^{2}+y^{2}+z^{2}+h^{2}\kappa^{2})(\pi^{2}-z^{2})^{2}}\left[\frac{16x^{2}z^{2}}{\pi^{2}h}\left[\sum_{n=1}^{\infty}\frac{h^{2}\lambda_{x}^{2}\sin\left(\frac{\pi n}{2}\right)}{n(\lambda_{x}^{2}x^{2}-4\pi^{2}n^{2}h^{2})}\right]^{2}+\frac{h^{3}\pi^{2}}{y^{2}}\right]\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}z$
@@
As long as the screening contribution in the sum in Eq. (18) is negligible ($\lambda_{x}\kappa\lesssim 1$), then the electrostatic energy $F_{\rho}$ grows with $\lambda_{x}$.
The nature of the domain walls may be complex due to the twist in the polarization $\mathbf{P}$. Experiments indicate that the domain wall may consist of surface disclination lines, with complex elastic distortions in the bulk near the disclinations [34, 35]. In addition, these walls will tend to have vanishing divergence of the polarization $\mathbf{P}$, so we expect no bound charges at the wall locations [4, 36]. An estimate for the energy of the wall (per unit length) due solely to disclinations would be $f^{\mathrm{disc}}_{\pi}\approx 2K$, with $K$ an elastic constant. Another possibility is that the domain wall is a solitonic, Bloch wall structure with an energy governed by the anchoring conditions, which we may call a $\pi$-wall/soliton as $\mathbf{P}$ flips by 180 across the wall [7]. The energy per unit length of one such $\pi$-soliton is given by $f^{\mathrm{soliton}}_{\pi}\approx 2\sqrt{2KhW}$, where $K$ is an elastic constant, $W$ is the anchoring strength, and $h$ is the cell thickness. A third possibility is that the energy of the wall is proportional to the cell thickness, which may occur if the elastic distortion of the wall extends across the entire thickness $h$, which may occur for thin cells, so that $f_{\pi}^{\mathrm{wall}}\approx Kh/\lambda_{\mathrm{wall}}$, where $\lambda_{\mathrm{wall}}$ is a characteristic width of the wall. We assume here that the characteristic width is independent of $h$. We will call this case the fixed-width domain wall. In summary, the total elastic cost of domain walls is approximately $F_{\mathrm{dw}}\approx 2f_{\pi}L^{2}/\lambda_{x}$, with $f_{\pi}$ either the disclination line pair $(f_{\pi}=f_{\pi}^{\mathrm{disc}})$, $\pi$-soliton ($f_{\pi}=f^{\mathrm{soliton}}_{\pi}$), or the fixed-width wall ($f_{\pi}=f_{\pi}^{\mathrm{wall}}$) energy density. More complicated possibilities or combinations of these three cases are also possible, but we will focus on just these three cases for simplicity.
@@
$\frac{F}{F_{0}}$ $=\frac{1}{4}+\frac{\kappa\lambda_{x}}{\pi^{2}}\sum_{n=1}^{\infty}\,\frac{[1-(-1)^{n}]}{n^{2}\sqrt{(2\pi n)^{2}+(\kappa\lambda_{x})^{2}}}+\,\frac{(\kappa\lambda_{\mathrm{dw}})^{2}L}{\kappa\lambda_{x}h},$ (19)
where $F_{0}=hP_{0}^{2}L(\epsilon\epsilon_{0}\kappa)^{-1}$ is a characteristic free energy and $\lambda_{\mathrm{dw}}=(2f_{\pi}\epsilon\epsilon_{0})^{1/2}P_{0}^{-1}$ is a characteristic length associated with the domain wall energy density $f_{\pi}$. A similar length was derived some time ago by balancing analogous energetic contributions [37]. We have made an additional assumption that we can take $h\kappa\gg 1$ in the left-over integration in Eq. (18), which will not change the location of the minimum of $F$ with respect to $\lambda_{x}$. We plot the dimensionless free energy $F/F_{0}$ in Eq. (19) (using the first 1000 non-zero terms in the sum) versus $\kappa\lambda_{x}$ in Fig. 6b for various values of $\kappa\lambda_{\mathrm{dw}}\sqrt{L/h}$, with $L$ the linear extent of the sample. Numerically minimizing the function in Eq. (19) demonstrates that for $\kappa\lambda_{\mathrm{dw}}\sqrt{L/h}<0.62$, the free energy curve has a global minimum at a finite value of $\lambda_{x}$. For larger values, the free energy minimum corresponds to $\lambda_{x}\rightarrow\infty$ and a uniform polarization state. In the low screening limit $\kappa\lambda_{\mathrm{dw}}\ll\sqrt{h/L}$, the preferred wavelength is at
@@
Model of pie-slice domains
Figure 7: Pie-slice domain model. a Schematic of the polarization orientation near the cell surface for a cell pre-patterned with a $+1$ defect. Apart from the domain walls (along the $x$ and $y$ axes and along the dashed line), the polarization $\mathbf{P}$ runs along the radial direction $\hat{\mathbf{r}}$. The polarization configuration has $n_{\theta}=4$ pairs of pie-slice sectors of anti-parallel $\mathbf{P}$. b Plot of the rescaled free energy $F/F_{0}$ in Eq. (26) of a pie-slice pattern with $n_{\theta}$ cuts for various values of $h\kappa$, with $\kappa$ the inverse Debye screening length and $h$ the cell thickness. We fix $R/h=100$ and $\eta_{\theta}=30$. We see that the Coulomb and elastic energies balance to generate an optimal value of $n_{\theta}$ (the free energy minimum). The dashed vertical line indicates the optimal value, $\eta_{\theta}$, in the low screening $\kappa\rightarrow 0$ limit.
We now consider the $+1$ aster defect anchoring condition, with $\mathbf{n}_{0}=\hat{\mathbf{r}}$ the preferred orientation at the top and bottom cell surfaces. A pure $+1$ aster defect in the polarization vector, $\mathbf{P}=P_{0}\hat{\mathbf{r}}$, has a corresponding bound charge density distribution
@@
The distribution of bound charge at the defect is energetically costly, so the $\mathrm{N}_{\mathrm{F}}$ prefers to reorient along the $\hat{\bm{\theta}}$ direction to create a net neutral charge configuration, forming a series of $n_{\theta}$ pairs of alternating $\mathbf{P}$ domains (with $\mathbf{P}\parallel\pm\hat{\mathbf{r}}$), as shown in Fig. 7(a). We also expect a $\pi$-twist along the $z$-direction, but we will neglect the $z$-dependence for simplicity, focusing on the SDs (see Fig. 2).
Consider the anchoring-imposed splay polarization $\mathbf{P}$ configuration in a circular domain of radius $R$. Far from the defect, as $R\rightarrow\infty$, the polarization will be uniform and we expect to see striped domains with the preferred wavelength $\lambda_{x}^{*}$ calculated in the previous section. Near the defect, we expect the bound charge distribution in the bulk should play a more significant role. Assuming the polarization remains in the $xy$ plane and does not depend on the distance $r$ from the defect, the form of the polarization vector of such a configuration is